# Alignment without Over-optimization: Training-Free Solution for Diffusion Models

### v1

@@ Abstract
Diffusion models excel in generative tasks, but aligning them with specific objectives while maintaining their versatility remains challenging. Existing fine-tuning methods often suffer from reward over-optimization, while approximate guidance approaches fail to optimize target rewards effectively. Addressing these limitations, we propose a training-free sampling method based on Sequential Monte Carlo (SMC) to sample from the reward-aligned target distribution. Our approach, tailored for diffusion sampling and incorporating tempering techniques, achieves comparable or superior target rewards to fine-tuning methods while preserving diversity and cross-reward generalization. We demonstrate its effectiveness in single-reward optimization, multi-objective scenarios, and online black-box optimization. This work offers a robust solution for aligning diffusion models with diverse downstream objectives without compromising their general capabilities. Code is available at https://github.com/krafton-ai/DAS.
@@ 1 Introduction
To address these limitations, we propose Diffusion Alignment as Sampling (DAS), a training-free approach that both achieves effective reward alignment and preserves model generalization. To guide latents toward high-reward samples, DAS leverages multiple candidate latents through Sequential Monte Carlo (SMC) sampling, averaging out errors in estimated corrections to enable sampling from a reward-aligned target distribution. By carefully designing intermediate target distributions with tempering techniques, DAS achieves high sample efficiency with multiple candidates, as we demonstrate both theoretically and empirically.
@@
  • We propose DAS, a training-free method for aligning diffusion models with arbitrary rewards while preserving general capabilities.
@@ 2.1 Fine-tuning Diffusion Models for Alignment
Aligning pretrained models through fine-tuning has been extensively studied in language models (Ziegler et al., 2020; Ouyang et al., 2022; Rafailov et al., 2023). For diffusion models, several approaches have emerged. Lee et al. (2023) and Wu et al. (2023b) employ supervised fine-tuning with preference-based reward models. Black et al. (2023) and Fan et al. (2024) formulate sampling as a Markov decision process and apply reinforcement learning (RL) to maximize rewards. Xu et al. (2024); Clark et al. (2024) and Prabhudesai et al. (2024) fine-tune by direct backpropagation through differentiable reward models. These approaches, however, face challenges with reward over-optimization (Gao et al., 2022; Coste et al., 2024), which may distort alignment or reduce sample diversity. KL regularization has been proposed as a mitigation strategy (Fan et al., 2024; Uehara et al., 2024a), inspired by its success in language models (Stiennon et al., 2020; Ouyang et al., 2022; Korbak et al., 2022). Section 3.2 examines the limitations of this approach, focusing on the mode-seeking behavior observed in the context of variational inference. While diffusion-based samplers (Zhang & Chen, 2022; Vargas et al., 2023; Berner et al., 2024; Sanokowski et al., 2024) use similar training objective to sample from multimodal, unnormalized target density, the fine-tuning setup makes training more susceptible to mode collapse (Appendix E). Alternatively, Zhang et al. (2024) approached over-optimization in RL fine-tuning through inductive and primacy biases.
@@ 3.1 Problem Setup: Aligning Diffusion Models with Rewards
However, this approach may lead to reward over-optimization (Gao et al., 2022), disregarding the pre-trained distribution. To mitigate this, we employ KL regularization (Korbak et al., 2022; Uehara et al., 2024a):
@@
Before we continue, we introduce binary optimality variable $\mathcal{O}\in\{0,1\}$ with $p(\mathcal{O}=1|x)\propto\exp(r(x)/\alpha)$, where samples with high reward are interpreted as more likely ”optimal”. Then the posterior $p(x|\mathcal{O}=1)$ characterizes the distribution of samples that achieve high rewards. Using Bayes’ rule with prior $p=p_{\text{pre}}$ give $p(x|\mathcal{O}=1)\propto p(x)p(\mathcal{O}=1|x)=p_{\text{pre}}(x)\exp\left(r(x)/\alpha\right)\propto p_{\text{tar}}(x)$, revealing the equivalence between two perspectives. We drop ’=1’ from now on following common convention.
@@ 3.2 Limitations of Existing Methods
Figure 1: SMC method excels in sampling from the target distribution compared to existing approaches. Left of dashed line: Samples from pre-trained model trained on mixture of Gaussians, reward-aligned target distribution $p_{tar}$. Right of dashed line: methods for sampling from $p_{tar}$ including previous methods (RL, direct backpropagation, approximate guidance) and ours using SMC. Top: reward $r(X,Y)=-X^{2}/100-Y^{2}$, bottom: reward $r(X,Y)=-X^{2}-(Y-1)^{2}/10$. EMD denotes sample estimation of Earth Mover’s Distance, also known as Wasserstein distance between the sample distribution using each method and the target distribution. Our SMC-based method outperforms existing approaches in capturing multimodal target distributions, as evidenced by lower EMD and successful sampling from all modes. Note that samples may exist outside the grid.
Previous approaches to sampling from the target distribution (Equation 3) primarily fall into two categories: fine-tuning and direct sampling using approximate guidance. In this subsection, we first demonstrate how these approaches struggle to sample from multimodal target distributions, even for simple Gaussian mixtures, and explain their limitations leading to potential failures.
@@
We first investigate the source of the mode-seeking behavior of fine-tuning methods. Fine-tuning methods can be interpreted as variational inference, with the following objective (Rafailov et al. (2023) Appendix A.1) :
@@
$\nabla_{x_{t}}\log p(\mathcal{O}|x_{t})\approx\frac{1}{\alpha}\nabla_{x_{t}}\hat{r}(x_{t})$ (6)
where $\hat{r}(\cdot)\coloneqq r(\hat{x}_{0}(\cdot))$. However, predicting clean data with noisy data introduces errors, especially at the beginning of sampling (He et al., 2024) where the noise is large, making it difficult to sample exactly from $p_{\text{tar}}$.
@@ 3.3 Sampling from Reward-aligned Target Distribution via Tempered SMC
  • Sequence of intermediate target distributions $\pi_{t}(x_{t})=\tilde{\gamma}_{t}(x_{t})/\mathcal{Z}_{t}$ for $t=0:T$ that bridge between the prior $\pi_{T}$ and target distribution $\pi_{0}$, where $\gamma_{t}$ is unnormalized density of $\pi_{t}$
@@
    $\bar{\pi}_{t}(x_{t:T})\coloneqq\pi_{t}(x_{t})\prod_{s=t+1}^{T}L_{s}(x_{s}|x_{s-1})$ (7)
  • Proposals, or transition kernels $m_{t-1}(x_{t-1}|x_{t})$ for sequential sampling
@@
    $w_{t-1}(x_{t-1},x_{t})\coloneqq\frac{\bar{\pi}_{t-1}(x_{t-1:T})}{\bar{\pi}_{t}(x_{t:T})m_{t-1}(x_{t-1}|x_{t})}\propto\frac{\tilde{\gamma}_{t-1}(x_{t-1})}{\tilde{\gamma}_{t}(x_{t})}\frac{L_{t}(x_{t}|x_{t-1})}{m_{t-1}(x_{t-1}|x_{t})}$ (8)
    in which proposed particles are resampled from $\text{Multinomial}(x_{t-1}^{1:N};w_{t-1}^{1:N})$
@@ 3.3.2 Intermediate Targets: Approximate Posterior with Tempering
$\pi_{t}(x_{t})\propto p_{t}(x_{t})p(\mathcal{O}|\hat{x}_{0}(x_{t}))^{\lambda_{t}}\propto p_{t}(x_{t})\exp\left(\frac{\lambda_{t}}{\alpha}\hat{r}(x_{t})\right)\eqqcolon\tilde{\gamma}_{t}(x_{t})$ (11)
@@ 3.3.3 Proposal: Approximating Locally Optimal Proposal
$m_{t-1}^{*}(x_{t-1}|x_{t})\propto\exp\left(-\frac{1}{2\sigma_{t}^{2}}\|x_{t-1}-\mu_{\theta}(x_{t},t)\|^{2}+\frac{\lambda_{t-1}}{\alpha}\hat{r}(x_{t-1})\right)$ (12)
@@
$m_{t-1}(x_{t-1}|x_{t})\coloneqq\mathcal{N}\left(\mu_{\theta}(x_{t},t)+\sigma_{t}^{2}\frac{\lambda_{t-1}}{\alpha}\nabla_{x_{t}}\hat{r}(x_{t}),\sigma_{t}^{2}I\right)$ (13)
@@
$w_{t-1}(x_{t-1},x_{t})=\frac{\tilde{\gamma}_{t-1}(x_{t-1})}{\tilde{\gamma}_{t}(x_{t})}\frac{L_{t}(x_{t}|x_{t-1})}{m_{t-1}(x_{t-1}|x_{t})}=\frac{p_{\theta}(x_{t-1}|x_{t})\exp\left(\frac{\lambda_{t-1}}{\alpha}\hat{r}(x_{t-1})\right)}{m_{t-1}(x_{t-1}|x_{t})\exp\left(\frac{\lambda_{t}}{\alpha}\hat{r}(x_{t})\right)}$ (14)
@@ 4.1.2 Results
Preserving diversity while optimizing rewards. Figure 3 showcases samples generated from the prompt ”crocodile,” aimed at maximizing aesthetic score. Our approach demonstrates superior aesthetic appeal while preserving sample diversity and pre-trained features of the animal. In contrast, samples from fine-tuning methods deviate significantly from the pre-trained model’s output and exhibit less diversity in colors, backgrounds, and appearances, indicating reward over-optimization.
Improving T2I alignment. Notably, in Figure 2, DAS substantially outperforms fine-tuning methods for the PickScore task across all metrics. To check whether the quantitative results align with actual human preferences, Figure 4 visualizes samples targeted to maximize PickScore across six categories: color, count, composition, location, style, and generating unusual scenes. DAS successfully generates aligned images with high visual appeal, even compared to fine-tuning baselines, thus effectively aligning the samples with human preferences. We provide additional results in Appendix F .
@@
Ablation on tempering We additionally conducted an ablation study on tempering, using the aesthetic score as target reward and ImageReward as evaluation for T2I alignment and cross-reward generalization. We also tested various tempering strategies, including $\gamma=0.008$, $\gamma=0.024$ (where $\lambda_{t}$ reaches 1 after 90 or 30 steps, respectively), and adaptive tempering (A.2). As shown in Figure 5, without tempering, SMC suffers from over-optimization, even with 32 particles, resulting in low ImageReward. In contrast, with tempering, both high aesthetic scores and ImageReward are achieved using only 4 or 8 particles, greatly improving efficiency regardless of tempering schemes. Tempering also reduces deviation from the latent manifold (He et al., 2024), indicating fewer generations of OOD samples. These findings align with the theoretical predictions in Section 3.3.3 and 3.3.4. While the performance of DAS is robust to tempering schemes, we recommend low $\gamma$ with a tuned $\alpha$ for the best balance of quality and efficiency in general.
@@ 4.2 Multi Rewards
Figure 6: Comparison of multi-objective optimization. (a) DAS achieve Pareto-optimal trade-off between the two rewards. (b) Prompt: ”classic cars on a city street with people and a dog”. Interestingly, DAS consistently generates ’dog’ in the prompt when aesthetic score weight gets lower than 0.9, while baselines fail to generate the dog for some images.
@@ 4.3 Online Black-box Optimization
Non-differentiable rewards. Reward maximization often involves non-differentiable or computationally expensive models. While DAS requires differentiable rewards for guidance, it can handle general rewards by posing the reward as black-box reward and learning a differentiable surrogate model with online feedback. We further demonstrate this using JPEG compressibility in F.3.
@@ 5 Conclusions
We introduce DAS, a training-free method using Sequential Monte Carlo sampling to align diffusion models with rewards. DAS optimizes rewards while preserving generalization without fine-tuning. In single and multi-reward experiments, DAS achieves comparable or superior target rewards to fine-tuning methods while excelling in diversity and cross-reward generalization. The online optimization results demonstrate DAS’s ability to efficiently explore diverse, high-quality solutions. These findings establish DAS as a versatile and efficient approach for aligning diffusion models applicable to a wide range of objectives and scenarios while significantly reducing the cost and complexity of the alignment process.
@@
We have made several efforts to ensure the reproducibility of our work. We provide complete proofs for all theoretical results in Appendix C, including formal statements and proofs for Propositions 2 and 3 in Appendix C.2.4 and C.2.5. Detailed pseudocode for our full DAS algorithm is included in Appendix A, with versions with adaptive resampling (Algorithm 1), adaptive tempering (Algorithm 3) and adaptation to online setting (Algorithm 5). Appendix D contains comprehensive implementation details for our method and baselines, including hyperparameter settings, training and sampling procedures. We will release our full codebase upon publication to enable others to replicate our results, including implementations of DAS. We use publicly available datasets and evaluation metrics, with details of experiment setup provided in Section 4. Appendix F contains additional experimental results to supplement those in the main paper. By providing these materials, we aim to enable other researchers to reproduce our results and build upon our work. We are committed to addressing any questions or requests for additional information to further support reproducibility efforts.

### v2

# Test-time Alignment of Diffusion Models without Reward Over-optimization

@@ Abstract
Diffusion models excel in generative tasks, but aligning them with specific objectives while maintaining their versatility remains challenging. Existing fine-tuning methods often suffer from reward over-optimization, while approximate guidance approaches fail to optimize target rewards effectively. Addressing these limitations, we propose a training-free, test-time method based on Sequential Monte Carlo (SMC) to sample from the reward-aligned target distribution. Our approach, tailored for diffusion sampling and incorporating tempering techniques, achieves comparable or superior target rewards to fine-tuning methods while preserving diversity and cross-reward generalization. We demonstrate its effectiveness in single-reward optimization, multi-objective scenarios, and online black-box optimization. This work offers a robust solution for aligning diffusion models with diverse downstream objectives without compromising their general capabilities. Code is available at https://github.com/krafton-ai/DAS.
@@ 1 Introduction
To address these limitations, we propose Diffusion Alignment as Sampling (DAS), a test-time algorithm that both achieves effective reward alignment and preserves model generalization. Our approach leverages Sequential Monte Carlo (SMC) sampling (Del Moral et al., 2006) to enable high-reward diffusion guidance by evaluating multiple possible solutions in parallel and selecting the best candidates during the generation process. By carefully designing intermediate target distributions with tempering techniques, DAS achieves high sample efficiency in terms of required candidates, as we demonstrate both theoretically and empirically.
@@
  • We propose DAS, a test-time method for aligning diffusion models with arbitrary rewards while preserving general capabilities.
@@ 2.1 Fine-tuning Diffusion Models for Alignment
Aligning pretrained models through fine-tuning has been extensively studied in language models (Ziegler et al., 2020; Ouyang et al., 2022; Rafailov et al., 2023). For diffusion models, several approaches have emerged. Lee et al. (2023) and Wu et al. (2023b) employ supervised fine-tuning with preference-based reward models. Black et al. (2023) and Fan et al. (2024) formulate sampling as a Markov decision process and apply reinforcement learning (RL) to maximize rewards. Xu et al. (2024); Clark et al. (2024) and Prabhudesai et al. (2024) fine-tune by direct backpropagation through differentiable reward models. These approaches, however, face challenges with reward over-optimization (Gao et al., 2022; Coste et al., 2024), which may distort alignment or reduce sample diversity. KL regularization has been proposed as a mitigation strategy (Fan et al., 2024; Uehara et al., 2024a), inspired by its success in language models (Stiennon et al., 2020; Ouyang et al., 2022; Korbak et al., 2022). Section 3.2 examines the limitations of this approach, focusing on the mode-seeking behavior observed in the context of variational inference. While diffusion-based samplers (Zhang & Chen, 2022; Vargas et al., 2023; Berner et al., 2024; Sanokowski et al., 2024) use similar training objective to sample from multimodal, unnormalized target density, the fine-tuning setup makes training more susceptible to mode collapse (Appendix F). Alternatively, Zhang et al. (2024) approached over-optimization in RL fine-tuning through inductive and primacy biases.
@@ 3.1 Problem Setup: Aligning Diffusion Models with Rewards
However, this approach may lead to reward over-optimization (Gao et al., 2022), disregarding the pre-trained distribution. To mitigate this, we employ KL regularization (Uehara et al., 2024a):
@@
Before we continue, we introduce binary optimality variable $\mathcal{O}\in\{0,1\}$ with $p(\mathcal{O}=1|x)\propto\exp(r(x)/\alpha)$, where samples with high reward are interpreted as more likely optimal. Then the posterior $p(x|\mathcal{O}=1)$ characterizes the distribution of samples that achieve high rewards. Using Bayes’ rule with prior $p=p_{\text{pre}}$ give $p(x|\mathcal{O}=1)\propto p(x)p(\mathcal{O}=1|x)=p_{\text{pre}}(x)\exp\left(r(x)/\alpha\right)\propto p_{\text{tar}}(x)$, revealing the equivalence between two perspectives. We omit ‘=1’ hereafter following convention.
@@ 3.2 Limitations of Existing Methods
Figure 1: SMC method excels in sampling from the target distribution compared to existing approaches. Left of dashed line: Samples from pre-trained model trained on mixture of Gaussians, reward-aligned target distribution $p_{tar}$. Right of dashed line: methods for sampling from $p_{tar}$ including previous methods (RL, direct backpropagation, approximate guidance) and ours using SMC. Top: reward $r(X,Y)=-X^{2}/100-Y^{2}$, bottom: reward $r(X,Y)=-X^{2}-(Y-1)^{2}/10$. EMD denotes sample estimation of Earth Mover’s Distance. Our SMC-based method outperforms existing approaches in capturing multimodal target distributions, as evidenced by lower EMD and successful sampling from all modes. Note that samples may exist outside the grid.
Previous approaches to sampling from the target distribution (Equation 3) primarily fall into two categories: fine-tuning and direct sampling using approximate guidance. Here, we first demonstrate how these approaches struggle to sample from multimodal target distributions, even for simple Gaussian mixtures, and explain their limitations leading to potential failures.
@@
We first investigate the source of the mode-seeking behavior of fine-tuning methods. Fine-tuning methods can be seen as variational inference, with the following objective (Rafailov et al. (2023)):
@@
$\nabla_{x_{t}}\log p(\mathcal{O}|x_{t})\approx\frac{1}{\alpha}\nabla_{x_{t}}\hat{r}(x_{t}),$ (6)
where $\hat{r}(\cdot)\coloneqq r(\hat{x}_{0}(\cdot))$. However, predicting clean data with noise causes errors, especially early in sampling when the noise is large, making it difficult to sample exactly from $p_{\text{tar}}$ (He et al., 2024).
@@ 3.3 Sampling from Reward-aligned Target Distribution via Tempered SMC
  • Sequence of intermediate target distributions $\pi_{t}(x_{t})=\tilde{\gamma}_{t}(x_{t})/\mathcal{Z}_{t}$ for $t=0:T$ that bridge between the prior $\pi_{T}$ and target distribution $\pi_{0}$, where $\gamma_{t}$ is unnormalized density of $\pi_{t}$.
@@
    $\bar{\pi}_{t}(x_{t:T})\coloneqq\pi_{t}(x_{t})\prod_{s=t+1}^{T}L_{s}(x_{s}|x_{s-1}).$ (7)
  • Proposals, or transition kernels $m_{t-1}(x_{t-1}|x_{t})$ for sequential sampling.
@@
    $w_{t-1}(x_{t-1},x_{t})\coloneqq\frac{\bar{\pi}_{t-1}(x_{t-1:T})}{\bar{\pi}_{t}(x_{t:T})m_{t-1}(x_{t-1}|x_{t})}\propto\frac{\tilde{\gamma}_{t-1}(x_{t-1})}{\tilde{\gamma}_{t}(x_{t})}\frac{L_{t}(x_{t}|x_{t-1})}{m_{t-1}(x_{t-1}|x_{t})},$ (8)
    where proposed particles are resampled from $\text{Multinomial}(x_{t-1}^{1:N};w_{t-1}^{1:N})$.
@@ 3.3.2 Intermediate Targets: Approximate Posterior with Tempering
$\pi_{t}(x_{t})\propto p_{t}(x_{t})p(\mathcal{O}|\hat{x}_{0}(x_{t}))^{\lambda_{t}}\propto p_{t}(x_{t})\exp\left(\frac{\lambda_{t}}{\alpha}\hat{r}(x_{t})\right)\eqqcolon\tilde{\gamma}_{t}(x_{t}),$ (11)
@@ 3.3.3 Proposal: Approximating Locally Optimal Proposal
$m_{t-1}^{*}(x_{t-1}|x_{t})\propto\exp\left(-\frac{1}{2\sigma_{t}^{2}}\|x_{t-1}-\mu_{\theta}(x_{t},t)\|^{2}+\frac{\lambda_{t-1}}{\alpha}\hat{r}(x_{t-1})\right).$ (12)
@@
$m_{t-1}(x_{t-1}|x_{t})\coloneqq\mathcal{N}\left(\mu_{\theta}(x_{t},t)+\sigma_{t}^{2}\frac{\lambda_{t-1}}{\alpha}\nabla_{x_{t}}\hat{r}(x_{t}),\sigma_{t}^{2}I\right),$ (13)
@@
$w_{t-1}(x_{t-1},x_{t})=\frac{\tilde{\gamma}_{t-1}(x_{t-1})}{\tilde{\gamma}_{t}(x_{t})}\frac{L_{t}(x_{t}|x_{t-1})}{m_{t-1}(x_{t-1}|x_{t})}=\frac{p_{\theta}(x_{t-1}|x_{t})\exp\left(\frac{\lambda_{t-1}}{\alpha}\hat{r}(x_{t-1})\right)}{m_{t-1}(x_{t-1}|x_{t})\exp\left(\frac{\lambda_{t}}{\alpha}\hat{r}(x_{t})\right)},$ (14)
@@ 4.1.2 Results
Preserving diversity while optimizing rewards. Figure 3 showcases samples generated from the prompt “crocodile,” aimed at maximizing aesthetic score. Our approach demonstrates superior aesthetic appeal while preserving sample diversity and pre-trained features of the animal. In contrast, samples from fine-tuning methods deviate significantly from the pre-trained model’s output and exhibit less diversity in colors, backgrounds, and appearances, indicating reward over-optimization.
Improving T2I alignment. Notably, in Figure 2, DAS substantially outperforms fine-tuning methods for the PickScore task across all metrics. To check whether the quantitative results align with actual human preferences, Figure 4 visualizes samples targeted to maximize PickScore across six categories: color, count, composition, location, style, and generating unusual scenes. DAS successfully generates aligned images with high visual appeal, even compared to fine-tuning baselines, thus effectively aligning the samples with human preferences. We provide additional results in Appendix G .
@@
Ablation on design choices. We additionally conducted an ablation study on tempering, using the aesthetic score as target reward and ImageReward as evaluation for T2I alignment and cross-reward generalization. We also tested various tempering strategies, including $\gamma=0.008$, $\gamma=0.024$ (where $\lambda_{t}$ reaches 1 after 90 or 30 steps, respectively), and adaptive tempering (A.2). In Figure 5, without tempering, SMC suffers from over-optimization, even with 32 particles, resulting in low ImageReward. In contrast, with tempering, using only 4 or 8 particles can achieve both high aesthetic scores and ImageReward, greatly improving efficiency regardless of tempering schemes. Tempering also reduces deviation from the latent manifold (He et al., 2024), indicating fewer generations of OOD samples. These findings align with the theoretical predictions in Section 3.3.3 and 3.3.4. While the performance of DAS is robust to tempering schemes, we recommend low $\gamma$ with a tuned $\alpha$ for the best balance of quality and efficiency in general. Test-time scaling experiment in Appendix E further demonstrates that DAS can effectively leverage additional computation during inference.
@@ 4.2 Multi Rewards
Figure 6: Comparison of multi-objective optimization. (a) DAS achieves a Pareto-optimal trade-off between the two rewards. (b) Prompt: “classic cars on a city street with people and a dog”. Interestingly, DAS consistently generates ‘dog’ in the prompt when the aesthetic score weight gets lower than 0.9, while baselines fail to generate the dog for some images.
@@ 4.3 Online Black-box Optimization
Non-differentiable rewards. Reward maximization often involves non-differentiable or computationally expensive models. While DAS requires differentiable rewards for guidance, it can handle general rewards by posing the reward as black-box reward and learning a differentiable surrogate model with online feedback. We further demonstrate this using JPEG compressibility in G.3.
@@ 5 Conclusions
We introduce DAS, a test-time algorithm using Sequential Monte Carlo sampling to align diffusion models with rewards. DAS optimizes rewards while preserving generalization without fine-tuning. In single and multi-reward experiments, DAS achieves comparable or superior target rewards to fine-tuning methods while excelling in diversity and cross-reward generalization. The online optimization results demonstrate DAS’s ability to efficiently explore diverse, high-quality solutions. These findings establish DAS as a versatile and efficient approach for aligning diffusion models applicable to a wide range of objectives and scenarios while significantly reducing the cost and complexity of the alignment process.
@@
We have made several efforts to ensure the reproducibility of our work. We provide complete proofs for all theoretical results in Appendix C, including formal statements and proofs for Propositions 2 and 3 in Appendix C.2.4 and C.2.5. Detailed pseudocode for our full DAS algorithm is included in Appendix A, with versions with adaptive resampling (Algorithm 1), adaptive tempering (Algorithm 3) and adaptation to online setting (Algorithm 5). Appendix D contains comprehensive implementation details for our method and baselines, including hyperparameter settings, training and sampling procedures. We will release our full codebase upon publication to enable others to replicate our results, including implementations of DAS. We use publicly available datasets and evaluation metrics, with details of experiment setup provided in Section 4. Appendix G contains additional experimental results to supplement those in the main paper. By providing these materials, we aim to enable other researchers to reproduce our results and build upon our work. We are committed to addressing any questions or requests for additional information to further support reproducibility efforts.